OpenStudy (anonymous):

The length of a rectangle is three times its width. If the perimeter of the rectangle is 48ft, find its area

hartnn (hartnn):

so, L=3W perimeter = 2(L+W) = 48 can you find L and W from these 2 equations ?

OpenStudy (anonymous):

L=24-w w=24-L

hartnn (hartnn):

yes, put that in L=3w

OpenStudy (anonymous):

w=6? L=18?

hartnn (hartnn):

absolutely correct! :) good.

hartnn (hartnn):

area = l*w =... ?

OpenStudy (anonymous):

108 is the area?

hartnn (hartnn):

correct :)